OpenStudy (anonymous):

Find each percent of change and round to the nearest hundreth.=95 to76 And 36to 30 solve thnkx

OpenStudy (anonymous):

cool you here thanks*

OpenStudy (anonymous):

95-76/90 = 19/90

OpenStudy (anonymous):

19/90 = .2

OpenStudy (anonymous):

Thanks/*

OpenStudy (anonymous):

36-30/36= 6/36=0.16

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

ok